OpenStudy (anonymous):

expand: x+2 to the third power

OpenStudy (anonymous):

you mean (x+2)^3? if that is it, it's (x+2)(x+2)(x+2)=(x^2+4x+4)(x+2)=(x^3+2x^2+4x^2+8x+4x+8) simplifies to: \[x ^{3}+6x ^{2}+12x+8\]